# NNsight and NDIF: Democratizing Access to Open-Weight Foundation Model Internals

### v3

@@ Abstract
We introduce NNsight and NDIF, technologies that work in tandem to enable scientific study of very large neural networks. NNsight is an open-source system that extends PyTorch to introduce deferred remote execution. NDIF is a scalable inference service that executes NNsight requests, allowing users to share GPU resources and pretrained models. These technologies are enabled by the intervention graph, an architecture developed to decouple experiment design from model runtime. Together, this framework provides transparent and efficient access to the internals of deep neural networks such as very large language models (LLMs) without imposing the cost or complexity of hosting customized models individually. We conduct a quantitative survey of the machine learning literature that reveals a growing gap in the study of the internals of large-scale AI. We demonstrate the design and use of our framework to address this gap by enabling a range of research methods on huge models. Finally, we conduct benchmarks to compare performance with previous approaches. Code, documentation, and tutorials are available at https://nnsight.net/.
@@ 1 Introduction
NDIF (Figure 1c), an open-source cloud inference service that supports this idiom by behind-the-scenes user sharing of model instances to reduce the costs of large-scale AI research. Crucially, by executing intervention graphs from multiple users with safe co-tenancy, NDIF supports the expressiveness of NNsight while avoiding costly startup times and communication costs incurred by traditional high-performance computing (HPC). We measure the performance of NDIF against the traditional approach, and we also compare it with Petals (11), a peer-to-peer approach to reducing barriers to distributed computing with large AI models.
@@ 2 Surveying model availability and research usage
Compared to commercial APIs, models with openly downloadable parameters enable much more invasive experimentation and are a valuable resource for interpretability researchers (49; 10; 57; 31). To understand current research activity studying the internals of these open-weight models, we examine a set of 184 papers sourced from a recent survey of interpretability research (24) — see Appendix A for data curation details.
@@
Studying models with up to 70 billion parameters is possible for labs with access to high-end compute such as NVIDIA A100 or H100 80GB GPUs, but doing research beyond this scale represents a leap in technical complexity and resource requirements. For example, inference on a 530 billion parameter LLM uses 1TB of GPU memory, requiring many devices distributed across multiple nodes just to load model parameters in 16-bit precision (4; 20). Even if a scientist has access to such resources, they must integrate their experiment code with model sharding, parallelization, quantization, and other deployment specifics. Thus, despite evidence that many fascinating capabilities appear only in the largest models (15; 47; 60; 45; 50), studying them can seem unattainable.
@@ 3 A Framework for Experiments on Large-Scale AI
Research on the internals of pretrained neural networks follows one common pattern: Experimental code is inserted between computational steps of the network, interleaved with code previously written to run the network. Depending on the experiment, researchers’ code may inspect internal representations (9), perform interventions (58), collect gradients (54; 51; 5; 7; 33), modify parameters (38), or use and train the parameters of supplementary models (35; 63; 29; 56; 37).
Traditionally, researchers have organized their code by creating bespoke versions of neural network modifications and hooks for each experiment using callback systems (36; 59; 26, and see Figure 4a), or by designing custom neural network systems with built-in access points (40; 64). While tying experiments with network implementations is primarily a matter of code organization for small models, it burdens researchers with the responsibility and cost of model deployment and storage, increasing engineering challenges as parameter sizes increase. This pattern also makes it difficult for external research groups to reproduce large distributed neural network experiments without duplicating the environment.
@@
  • Expressiveness: The framework must allow the user to express all the experiments that they would have been able to express if they were using PyTorch on local hardware
@@ 3.1 Representing experiments as graphs
Our core innovation is a portable, serializable representation of an experiment, called the intervention graph, which dynamically captures modifications to the model’s computation graph. This graph can be stored in format, version-controlled, optimized, and sent to or retrieved from remote systems for execution. In this section, we formalize the notion of an intervention graph.
@@
Figure 3: Augmenting the computation graph $C$ (pink graph on the left) with an intervention graph (green, right). (a) A getter edge brings the output of a variable node $v_{1}$ in the computation graph to an apply node $a_{1}^{\prime}$ in the intervention graph. (b) A setter edge sends the output of $v_{2}^{\prime}$ in the intervention graph back to $a_{3}$ in the computation graph.
@@
Interventions are modifications to the computation graph $C$ that introduce additional nodes and edges. We define an intervention as $I=(V^{\prime},\,A^{\prime},\,E^{\prime},\,\mathcal{G},\,\mathcal{S}\}$, where $(V^{\prime},\,A^{\prime},\,E^{\prime})$ forms an intervention component $G$. Each $G$ inherits the properties of a computation graph. The sets $\mathcal{G}$ and $\mathcal{S}$, referred to as getters and setters, define how $G$ connects to $C$.
Integrating an intervention component into $C$ is called interleaving, and consists of two parts:
  • The getters $(\mathcal{G})$ connect variable nodes in $C$ to apply nodes in $G$, i.e., $\mathcal{G}\subseteq V\times A^{\prime}$.
  • The setters $(\mathcal{S})$ connect variable nodes in $G$ back to apply nodes in $C$, i.e., $\mathcal{S}\subseteq V^{\prime}\times A$.
The intervention graph of NNsight is the union of all individual intervention components, $\mathcal{I}=\bigcup_{1\leq i\leq k}G_{i}$. The final augmented computation graph $C^{\prime}$ is constructed by interleaving each intervention component with the original graph. $C^{\prime}$ is a new computation graph, inheriting the properties of $C$. When an intervention graph is sent to NDIF, the getters and setters are also included. Appendix B.1 discusses implementation details.
@@ 3.2 Familiar and expressive interventions
Because PyTorch code is used inside the tracing context, any experiment that can be written using standard PyTorch hooks can also be expressed using NNsight and will feel familiar to PyTorch users. NNsight wraps all 217 fundamental PyTorch tensor operations, supports built-in modules and custom neural network architectures, and includes infrastructure for loading models distributed by popular platforms such as HuggingFace. Users can express a wide range of interventions beyond simple probing and manipulation, including custom attention mechanisms, dynamic computation graph alterations, weight updates, and training loops, as well as computing gradients on backwards passes and gradient flow modifications.
@@ 3.3 Co-Tenancy and remote infrastructure
The NDIF service, in contrast to the NNsight system, is a multi-user runtime infrastructure designed to amortize costs and remove engineering burden from scientists by concurrently sharing inference. NDIF accepts intervention graphs created by NNsight, and executes them on large, preloaded models.
@@
Safe co-tenancy. NDIF assumes responsibility for ensuring that user experiments conducted on the same model and hardware do not interfere with each other by preventing direct access to model parameters and by enforcing a server-side whitelist of permitted operations that ensure complete separation of user experiments. The computation graph’s interleaving mechanism allows users to simulate parameter interventions without accessing or modifying the parameters themselves. NDIF also tracks and authorizes individual usage to maintain fair resource allocation.
@@
The code for creating NDIF infrastructure is freely available on GitHub, allowing users to create their own NDIF infrastructure for specialized use cases.
@@ 4 Performance and Evaluation
In this section, we evaluate the performance of NNsight when used remotely, both using exclusive allocations in the traditional high-performance computing setting, and using the shared NDIF inference service. We also compare to the Petals (11) distributed inference system, which can also reduce the costs of customized inference on huge models. Our evaluation focuses on the time required to load the model weights into memory, as well as the runtime of activation patching, a standard model intervention technique (58). For this purpose, we use a single batch of 32 examples from the Indirect Object Identification (IOI) dataset (59). We evaluate performance in three different settings: First, we measure the runtime when the models and interventions are executed entirely on a high-performance computing node (HPC). Then, we execute the models and interventions on the NDIF infrastructure, and contrast it against HPC execution times. Finally, we compare NDIF with etals (11), and evaluate their relative performance on standard inference tasks as well as on intervention tasks.
@@ 5 Related Work
Several other solutions for sharing hosted model resources have been developed, including VLLM (34), Huggingface TGI (19) and Nvidia TensorRT-LLM (41). These are systems for LLM inference acceleration that support large-scale multiuser model sharing. However, unlike NDIF, these systems only provide black-box API access to models, and do not permit model inspection or modification, which limits their utility for interpretability research. Also related to our work are systems such as S-LoRA (52), dLoRA (62), and Punica (18), which excel at efficient serving of many pretrained LoRA adapters in parallel. While these LoRA-focused systems enable one form of model modification, NNsight and NDIF provide finer-grained control that allows for a wider range of model interventions.
@@ 8 Reproducibility Statement
To promote transparency and ensure the reproducibility of our results, the complete NNsight and NDIF frameworks, along with comprehensive documentation and all relevant experiment code, will be made publicly available as open-source resources. This will allow researchers and practitioners to replicate our framework, related experiments and build upon our work.
@@ 9 Acknowledgements
The National Deep Inference Fabric and NNsight are supported by the U.S. National Science Foundation Award IIS-2408455. We are also grateful for the support of Open Philanthropy (ET, KP, SM, AS, NP, DB), Manifund Regrants (CR), and the Zuckerman Postdoctoral Fellowship (AM). We thank Brett Bode, William Gropp, Greg Bauer, and Volodymyr Kindratenko at NCSA for their help working with DeltaAI, which provides the primary computing infrastructure for NDIF, and we thank Thea Shaheen for invaluable assistance with infrastructure for the project at Northeastern. We thank the New England Research Cloud (NERC) for hosting the web frontend for NDIF. We also thank Callum McDougall for developing the NNsight tutorial in the ARENA 3.0 materials and Dmitrii Troitskii for running outreach and tutorial efforts. We thank Clément Dumas and Gonçalo Paulo for valuable feedback as early research adopters, Zhengxuan Wu for implementing pyvene on NNsight, and Zhongbo Zhu for valuable advice on serving infrastructure. Finally, we are grateful for the ongoing feedback and engagement of our community of users on Discord.
@@ Appendix B System Design
In this section, we provide additional details describing the implementation and design of NNsight and NDIF. Figure 8 provides an overview of the different components of the system, their responsibilities, and how they interact with one another.
Figure 8: An overview of the NNsight and NDIF remote system. (1) Researchers write experiment code using the NNsight API which is converted to an Intervention Graph. (2) The graph is serialized to a custom JSON format and sent as a request to the NDIF frontend server. The NDIF backend can host multiple model instances, each on a dedicated set of GPU nodes. For very large models, such as Llama 3.1 405B (21), model weights are distributed across many shards using tensor parallelism. (3) The router transfers the request to the head node (shard 0) of the requested model, via the Ray GCS Service (39; 48). (4) Shard 0 sends the request to all other shards of the model where it is then deserialized and executed. Each shard receives the full intervention graph, but only manages a slice of the model parameters. The Torch NCCL Head manages distributed model execution across allocated shards. (5) After the intervention graph has been executed, results are gathered at shard 0 and sent to the object store in the NDIF frontend. The shard 0 WebSocket client informs the WebSocket client on the local workstation about the completion of the intervention. (6) As soon as the local WebSocket client notes the intervention is complete it pulls the final results from the Object Store and inserts the result back into the local intervention graph. The research code can pull results from the intervention graph that are requested via .save().
@@ 27 Args:
Code Example 5: Using NNsight to train a LORA with remote execution, the Session context, and iterative interventions. First, the LORA class is defined that takes an Envoy as input. Then, the session context is used for remote training with an optimizer.

### v4

@@ Abstract
We introduce NNsight and NDIF, technologies that work in tandem to enable scientific study of very large neural networks. NNsight is an open-source system that extends PyTorch to introduce deferred remote execution. The National Deep Inference Fabric (NDIF) is a scalable inference service that executes NNsight requests, allowing users to share GPU resources and pretrained models. These technologies are enabled by the intervention graph, an architecture developed to decouple experimental design from model runtime. Together, this framework provides transparent and efficient access to the internals of deep neural networks such as very large language models (LLMs) without imposing the cost or complexity of hosting customized models individually. We conduct a quantitative survey of the machine learning literature that reveals a growing gap in the study of the internals of large-scale AI. We demonstrate the design and use of our framework to address this gap by enabling a range of research methods on huge models. Finally, we conduct benchmarks to compare performance with previous approaches. Code, documentation, and tutorials are available at https://nnsight.net/.
@@ 1 Introduction
NDIF (Figure 1c), an open-source cloud inference service that supports this idiom by providing behind-the-scenes user sharing of model instances to reduce the costs of large-scale AI research. Crucially, by executing intervention graphs from multiple users with safe co-tenancy, NDIF supports the expressiveness of NNsight while avoiding costly startup times and communication costs incurred by traditional high-performance computing (HPC). We measure the performance of NDIF against the traditional approach, and we also compare it with Petals (11), a peer-to-peer approach to reducing barriers to distributed computing with large AI models.
@@ 2 Surveying model availability and research usage
Compared to commercial APIs, models with openly downloadable parameters enable much more invasive experimentation and are a valuable resource for interpretability researchers (49; 10; 59; 31). To understand current research activity studying the internals of these open-weight models, we examine a set of 184 papers sourced from a recent survey of interpretability research (24) — see Appendix A for data curation details.
@@
Studying models with up to 70 billion parameters is possible for labs with access to high-end compute such as NVIDIA A100 or H100 80GB GPUs, but doing research beyond this scale represents a leap in technical complexity and resource requirements. For example, inference on a 530 billion parameter LLM uses 1TB of GPU memory, requiring many devices distributed across multiple nodes just to load model parameters in 16-bit precision (4; 20). Even if a scientist has access to such resources, they must integrate their experiment code with model sharding, parallelization, quantization, and other deployment specifics. Thus, despite evidence that many fascinating capabilities appear only in the largest models (15; 47; 62; 45; 50), studying them can seem unattainable.
@@ 3 A Framework for Experiments on Large-Scale AI
Research on the internals of pretrained neural networks follows one common pattern: Experimental code is inserted between computational steps of the network, interleaved with code previously written to run the network. Depending on the experiment, researchers’ code may inspect internal representations (9), perform interventions (60), collect gradients (54; 51; 5; 7; 33), modify parameters (38), or use and train the parameters of supplementary models (35; 65; 29; 58; 37).
Traditionally, researchers have organized their code by creating bespoke versions of neural network modifications and hooks for each experiment using callback systems (36; 61; 26, and see Figure 3a), or by designing custom neural network systems with built-in access points (40; 66). While tying experiments with network implementations is primarily a matter of code organization for small models, it burdens researchers with the responsibility and cost of model deployment and storage, increasing engineering challenges as parameter sizes increase. This pattern also makes it difficult for external research groups to reproduce large distributed neural network experiments without duplicating the environment.
@@
  • Expressiveness: The framework must allow the user to express all the experiments that they would have been able to express if they were using PyTorch on local hardware.
@@ 3.1 Representing experiments as graphs
Our core innovation is a portable, serializable representation of an experiment, called the intervention graph, which dynamically captures modifications to the model’s computation graph. This graph can be stored in JSON format, version-controlled, optimized, and sent to or retrieved from remote systems for execution. In this section, we formalize the notion of an intervention graph.
@@
Interventions are modifications to the computation graph $C$ that introduce additional nodes and edges. We define an intervention component $C^{\prime}=(V^{\prime},A^{\prime},E^{\prime})$, where $C^{\prime}$ is a computation graph representing a single user-specified edit. We define an intervention as $I=\{C^{\prime},\,\mathcal{G},\,\mathcal{S}\}$, where sets $\mathcal{G}$ and $\mathcal{S}$, referred to as getters and setters, determine how $C^{\prime}$ connects to $C$. Integrating an intervention component into $C$ is called interleaving, and consists of two parts:
  • The getters $(\mathcal{G})$ connect variable nodes in $C$ to apply nodes in $C^{\prime}$, i.e., $\mathcal{G}\subseteq V\times A^{\prime}$.
  • The setters $(\mathcal{S})$ connect variable nodes in $C^{\prime}$ back to apply nodes in $C$, i.e., $\mathcal{S}\subseteq V^{\prime}\times A$.
The intervention graph of NNsight is the union of all individual intervention components, $\mathcal{I}=\bigcup_{1\leq i\leq k}{C_{i}^{\prime}}$. The final augmented computation graph is constructed by interleaving each intervention component with the original graph, and represents a complete user experiment. Appendix B.1 discusses implementation details of how intervention graphs are constructed, and Figure 10 (Appendix E) shows an example of an abstract intervention graph.
@@ 3.2 Familiar and expressive interventions
Because PyTorch code is used inside the tracing context, any experiment that can be written using standard PyTorch hooks can also be expressed using NNsight and will feel familiar to PyTorch users. NNsight wraps all 217 fundamental PyTorch tensor operations, supports built-in modules and custom neural network architectures, and includes infrastructure for loading models distributed by popular platforms such as HuggingFace. Users can express a wide range of interventions beyond simple probing and manipulation, including custom attention mechanisms, dynamic computation graph alterations, weight updates, and training loops (see Code Example 5), as well as computing gradients on backwards passes and gradient flow modifications.
@@ 3.3 Co-Tenancy and remote infrastructure
Figure 4: An overview of the NNsight and NDIF remote system. Researchers write experiment code using the NNsight API which is converted to an Intervention Graph. The graph is serialized to a custom JSON format and sent as a request to the NDIF frontend server. The NDIF backend can host multiple model instances, each on a dedicated set of GPU nodes. For very large models, such as Llama 3.1 405B (21), model weights are distributed across many shards using tensor parallelism. The router transfers the request to the head node (shard 0) of the requested model, via the Ray GCS Service (39; 48). Shard 0 sends the request to all other shards of the model where it is then deserialized and executed. Each shard receives the full intervention graph, but only manages a slice of the model parameters. The Torch NCCL Head manages distributed model execution across allocated shards. After the intervention graph has been executed, results are gathered at shard 0 and sent to the object store in the NDIF frontend. The shard 0 WebSocket client informs the WebSocket client on the local workstation about the completion of the intervention. As soon as the local WebSocket client notes the intervention is complete it pulls the final results from the Object Store and inserts the result back into the local intervention graph. The research code can pull results from the intervention graph that are requested via .save().
The NDIF service, in contrast to the NNsight system, is a multi-user runtime infrastructure designed to amortize costs and remove engineering burden from scientists by concurrently sharing inference. NDIF accepts intervention graphs created by NNsight, and executes them on large, preloaded models. In Figure 4, we provide an overview of the different components of the system, their responsibilities, and how they interact with one another.
@@
Safe co-tenancy. NDIF assumes responsibility for ensuring the safe and fair use of its shared model resources via multiple protocols. Interference between user experiments is prevented by the intervention graph’s interleaving mechanism, which enables users to simulate interventions without permanently modifying the model’s original computation graph. Model parameters cannot be directly accessed; instead, users interact with a copy of the parameters. Finally, users can only access models hosted on NDIF if they have been authorized by the model providers. This restriction is naturally enforced, as submitting a request to NDIF through the NNsight API requires downloading the model’s meta version from Hugging Face, where access is managed directly by the model providers.
@@
The code for creating NDIF infrastructure is freely available on GitHub at https://github.com/ndif-team/ndif, allowing users to create their own NDIF infrastructure for specialized use cases.
@@ 4 Performance and Evaluation
In this section, we evaluate the performance of NNsight when used remotely, both using exclusive allocations in the traditional high-performance computing setting, and using the shared NDIF inference service. We also compare to the Petals (11) distributed inference system, which can also reduce the costs of customized inference on huge models. Our evaluation focuses on the time required to load the model weights into memory, as well as the runtime of activation patching, a standard model intervention technique (60). For this purpose, we use a single batch of 32 examples from the Indirect Object Identification (IOI) dataset (61). We evaluate performance in three different settings: First, we measure the runtime when the models and interventions are executed entirely on a high-performance computing node (HPC). Then, we execute the models and interventions on the NDIF infrastructure, and contrast it against HPC execution times. Finally, we compare NDIF with Petals (11), and evaluate their relative performance on standard inference tasks as well as on intervention tasks.
@@ 5 Related Work
Several other solutions for sharing hosted model resources have been developed, including VLLM (34), Huggingface TGI (19) and Nvidia TensorRT-LLM (41). These are systems for LLM inference acceleration that support large-scale multiuser model sharing. However, unlike NDIF, these systems only provide black-box API access to models, and do not permit model inspection or modification, which limits their utility for interpretability research. Also related to our work are systems such as S-LoRA (52), dLoRA (64), and Punica (18), which excel at efficient serving of many pretrained LoRA adapters in parallel, and systems such as FusionAI and FusionLLM (55; 56) which support decentralized training. While these LoRA-focused systems enable one form of model modification, NNsight and NDIF provide finer-grained control that allows for a wider range of model interventions.
@@ 8 Reproducibility Statement
To promote transparency and ensure the reproducibility of our results, the complete NNsight and NDIF frameworks, along with comprehensive documentation and all relevant experiment code, are available at https://github.com/ndif-team/nnsight and https://github.com/ndif-team/ndif, respectively. We hope this will allow researchers and practitioners to build upon our work.
@@ 9 Acknowledgements
The National Deep Inference Fabric (NDIF) and NNsight are supported by the U.S. National Science Foundation Award IIS-2408455. We are also grateful for the support of Open Philanthropy (ET, KP, SM, AS, NP, DB), Manifund Regrants (CR), and the Zuckerman Postdoctoral Fellowship (AM). We thank Brett Bode, William Gropp, Greg Bauer, and Volodymyr Kindratenko at NCSA for their help working with DeltaAI, which provides the primary computing infrastructure for NDIF, and we thank Thea Shaheen for invaluable assistance with infrastructure for the project at Northeastern. We thank the New England Research Cloud (NERC) for hosting the web frontend for NDIF. We also thank Callum McDougall for developing the NNsight tutorial in the ARENA 3.0 materials. We thank Clément Dumas and Gonçalo Paulo for valuable feedback as early research adopters, Zhengxuan Wu for implementing pyvene on NNsight, and Zhongbo Zhu for valuable advice on serving infrastructure. We also acknowledge the many early testers for the Llama 3.1 405B pilot program. Finally, we are grateful for the ongoing feedback and engagement of our community of users on Discord.
@@ Appendix B System Design
In this section, we provide additional details describing the implementation and design of NNsight and NDIF.
@@ 27 Args:
Code Example 5: Using NNsight to train a LORA with remote execution, the Session context, and iterative interventions. NNsight supports creating parameters and optimizers remotely. Using ‘model.trace‘ as a forward pass, a dataset can be iterated over to optimize the remote parameters. The normal workflow of calculating a loss, executing a backwards pass, and taking optimizations steps is expressed just as if this were code executing on the user’s local workstation. After executing the training loop, the optimized parameters are sent back to the user’s workstation to be used locally, or to be sent back to the remote server for future applications of parameters.
@@ Appendix E Equivalence of Theano’s Many-to-Many and NNsight’s Many-to-One Apply Nodes
Figure 10 shows an example of how getters and setters interact with an intervention graphs variable and apply nodes on an abstract computation graph.
Figure 10: Augmenting the computation graph $C$ (pink graph on the left) with an intervention graph (green, right). (a) A getter edge brings the output of a variable node $v_{1}$ from the computation graph to an apply node $a_{1}^{\prime}$ in the intervention graph. (b) A setter edge sends the output of $v_{2}^{\prime}$ in the intervention graph back to $a_{3}$ in the computation graph.